[figure: 9.7 cm]

JAHN-TELLER EFFECT AND SUPEREXCHANGE IN HALF-DOPED MANGANITES

Olga Sikora and Andrzej M. Oleś

Institute of Physics, Jagellonian University, Reymonta 4, 30-059 Kraków, Poland

While the A-type antiferromagnetic (AF) phase in the undoped LaMnO_{3} is well understood and follows from the superexchange interactions due to e_{g} and t_{2g} electron excitations [1], the stability of the charge exchange (CE) phase in half doped manganites remains one of the outstanding problems. We investigate the stability of this phase within a microscopic model which describes a single plane as in La_{0.5}Sr_{1.5}MnO_{4} compound. It includes: the kinetic energy ∝ t of e_{g} electrons moving in nearly degenerate orbitals and interacting by local Coulomb interaction U, the Jahn-Teller (JT) term ∝ E_{JT} which induces orbital polarization, and the superexchange interactions ∝ J, as derived for the pairs of Mn^{3+}–Mn^{4+}, Mn^{3+}–Mn^{3+}, and Mn^{4+}–Mn^{4+} ions [2]. Treating the electron-electron interactions within the Hartree approximation, and the superexchange within the mean-field approximation, we determine the magnetic, orbital and charge order parameters in various magnetic phases: G-type AF, A-type AF, C-type AF, and CE phase as functions of E_{JT} and J. By investigating the phase diagram at T = 0 we conclude that the cooperative JT effect plays a prominent role in stabilizing the CE phase, while its magnetic order could be explained only by somewhat stronger superexchange than that which follows from the spectroscopic parameters.

[1] L.F. Feiner and A.M. Oleś, Phys. Rev. B 59, 3295 (1999).

[2] A.M. Oleś and L.F. Feiner, Phys. Rev. B 65, 052414 (2002).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Olga Sikora

Address for correspondence : Institute of Physics Jagellonian University Reymonta 4 PL-30059 Kraków

Email address : sikora@alphetna.if.uj.edu.pl